I didn't know where to put this question but the equation is 3^(2x+1)=1. The answer is apparently -1/2 but I have no idea how to get this.
Answer by lwsshak3(11628)

I didn't know where to put this question but the equation is
3^(2x+1)=1
take log of both sides
(2x+1)log(3)=log(1)
(2x+1)log(3)=0
2x+1=0
x=-1/2